OpenStudy (anonymous):

One circle has a circumference of 12cm. Another circle has a circumference of 32cm. What is the radius of the smaller circle to the radius of the larger circle? A. 3:8 B. 1:3 C. 1:3 D. 3:1

OpenStudy (anonymous):

I came up with 3.2

OpenStudy (anonymous):

12/3.14=3.8/2=1.9 32/3.14=10.2/2=5.1 got it the answer is 3.1

OpenStudy (anonymous):

3.1 is not right